Shift report 20010805 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Aug 05 2001 - 16:21:13 EDT

  • Next message: Apache: "Shift report 20010806 16:00-24:00"

    ------------------------------------------------------------
    Date: Sun Aug  5 16:21:12 2001, Shift: 08:00-16:00
    Supervisor: PC, with: CR
    Log book#7 pages 24 - 28
    
    Shift Summary:
    ------------------------------------------------------------
    1) End of run 3994 taken over from night shift.
    Nothing more.
    
    2) MCR switched polarity on magnets after and agreed with 
    next shift crew to have :
    D5 - 2kGB (350 Amp)
    D1-D4 (1/5 field B) to continue run plan. 
    The alternative was to use 1/3 B. 
    
    3) With the polarity switched we turned all fields on the
    pet page to B. But we realised when the next shift crew came
    that the values in the Amps M field was either extremely 
    high or extremely low for all settings. We then switched the
    mode from B to A and everything was almost normal again.
    Magnet D3 have status 0x408 and the other magnets have
    status 0x8 and the DoIt field is Red when pushed, but
    the Amps come on. We left a phone message for Dana.
    
    The T4 LED panel is still in FAULT. Means anything ?
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Aug 05 2001 - 16:21:45 EDT